OpenStudy (anonymous):

Standard Form Help. 1. Determine the equation of a line in standard form that represents Jake’s training progress. His progress corresponds to the points (1, 9.50) and (4, 6.75).

OpenStudy (anonymous):

(1, 9.50) and (4, 6.75) with those two points, what did you get for the slope ?

OpenStudy (anonymous):

ok, now we can do y - 9.5 = -11/12(x - 1) multiply both sides by 12 to get 12y - 12*9.5 = -11(x-1) or 12y - 114= -11x + 11 add 11x to both sides 11x +12y -114= 11 add 114 to both sides 11x +12y = 125

OpenStudy (anonymous):

it is -0.91667 or -11/12

OpenStudy (anonymous):

medal plz, if helped

OpenStudy (anonymous):

2. Determine the equation of a line in standard form that represents Samantha’s training progress. Her progress corresponds to the points (1, 8.75) and (5, 5.75)

OpenStudy (anonymous):

I think I figured out the answer, but could you show the steps?